A bullet leaves the rifilsof mass one kg and the recoils therby with the velocity of 30 cm/s if mass of bullet is 3 g find the velocity of bullet

Grade:11

1 Answers

Arun
6 years ago
Dear Tejas
 
From conservation of momentum
P1 = P2
m1 * V1 = m2 * V2
1 * 30 = 0.003 * V2
V2 = 10000 cm/sec
 
Regards
Arun (askIITians forum expert)